OpenStudy (anonymous):

solve integral of sinx^2

OpenStudy (anonymous):

I don't think we can solve integral of \[\sin {x^2}\]. You can see here : http://www.wolframalpha.com/input/?i=%5Cint+%7Bsin+x%5E2%7D

OpenStudy (anonymous):

can someone show steps ?

OpenStudy (anonymous):

It means we cannot find the solution.

OpenStudy (anonymous):

ah, thanks, i wasn't sure

OpenStudy (anonymous):

I think we can find it

OpenStudy (anonymous):

Can you show it here?

OpenStudy (anonymous):

Integral of sin (f(x)) is -cos(f(x))/f'(x) I think

OpenStudy (anonymous):

:) using your above formula, I have the result is \[f(x)=-\frac{\cos x^2}{2x}+C\], you try to find the derivation of f(x), if it equals \[\sin x^2\], you can deduce your formula is true. You see here : http://mathforum.org/library/drmath/view/51877.html

OpenStudy (anonymous):

you have a confusion, the integral is \[\int\limits {\sin x^2}dx\], not \[\int\limits\limits {\sin^2 x}dx\]